OpenStudy (anonymous):

What is the radius of a circle in which a 30° arc is 2 pi inches long

OpenStudy (anonymous):

30/360 = 2pi/circumference 1/12(circumference) = 2pi circumference = 2pi/(1/12) circumference = 24pi circumference = 2r(pi) 24pi = 2r(pi) 24 = 2r r = 12 in.